Shift report 20000719 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Jul 19 2000 - 08:21:44 EDT

  • Next message: Flemming Videbaek: "Fw: operator account setup changed!"

    ------------------------------------------------------------
    Date: Wed Jul 19 08:21:43 2000, Shift: 00:00-08:00
    Supervisor: Michael, with: none
    Log book#3 pages 28 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    The beam was down most of the shift  due to power surply
    problems in the booster. We had about 15 minutes of mainly
    blue beam. There were some DAQ problems. Supermon seems
    to have reverted to an old version after the software 
    changes of yesterday evening. Tried to get the BB TDC
    calibration done but BB Map shows and error in DAQ.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 19 2000 - 08:58:58 EDT